# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 6.2.2. Security of Containerized Clouds
Nicolae et al. (Casalicchio, 2019) analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework. Their research identifies potential attack scenarios and security enforcement mechanisms, aiming to enhance security guarantees for cloud operators in a multi-cloud setting. Mohammad et al. (Bokhari et al., 2018) offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models. They discuss the security requirements for public and private clouds across different service models, aiming to provide researchers with a comprehensive understanding of cloud computing’s potential and security challenges. Sari et al. (Sultan et al., 2019) survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape. Their analysis covers both software-based solutions utilizing Linux kernel features and hardware-based solutions for enhanced security. The study aims to clarify container security requirements and encourage further research in addressing potential vulnerabilities and attacks. Hendrik et al. (Graupner et al., 2015a) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective. They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control. Rohan et al. (Gupta et al., 2016) propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption. This approach aims to secure data while optimizing resource usage and cost, highlighting the benefits of container technology in improving cloud computing security and efficiency.
@@ 6.3. Conclusive Summary
Current survey-based studies and systematic reviews, discussed above (e.g., (Casalicchio and Iannucci, 2020) (Pahl et al., 2017)), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems. These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends.
Our study distinguishes itself from prior research in several ways. First, our work systematically categorizes container utilization in multi-cloud environments into coherent categories and themes. Some of the key themes are ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided. Furthermore, we explored the QAs and associated tactics for containerization in multi-cloud environment. This contribution offers practitioners actionable insights into tailoring their containerization strategies to meet desired quality standards. Lastly, our study introduces four distinct frameworks consisting of challenges and solutions related to security, automation, deployment, and monitoring within the context of containerization in multi-cloud environment.
@@ 7. Conclusions
This SMS presents the current state of research regarding containerization in multi-cloud environment, focusing on aims of selected studies, the roles of containers, containers implementation strategies, architectural patterns and strategies, quality attributes and corresponding tactics. Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment. From an initial pool of extensive search results, 86 relevant studies are were finally chosen to answer the research questions presented in Table 1. The key findings of this SMS are listed below.
  • The findings of this SMS reveal various insights regarding SMS demographics, publishers, publication types, authors’ affiliations, and research themes. The yearly distribution reached its peak in 2016, signifying a shifting level of interest over time. The field of publishing is largely dominated by IEEE (72.09%), with conferences emerging as the primary avenue for publication. Notably, academia (93.02%) stands out as the leading affiliation among authors. The prominent research themes encompass Orchestration and Management, Scalability and High Availability, Performance and Optimization, and Security and Privacy. These themes underscore crucial aspects such as container management, scalability, performance, and security within the context of multi-cloud environment.
  • This SMS identifies and classifies a total of 74 patterns and strategies across 10 subcategories and 4 categories for container-based applications in multi-cloud environment. These patterns and strategies encompass a wide range of aspects such as cloud architecture models, communication and networking, deployment, service models, cloud management, container management, edge computing, IoT strategies, security, resilience, fault-tolerance, and cloud migration. These findings are significant for the application and dvancement of containerization-based applications in multi-cloud environment, providing valuable insights into optimizing architecture, communication, management, security, and migration strategies.
@@
  • This SMS presents a framework for the automation challenge-solution framework for containerized applications in multi-cloud environment. The automation challenge-solution framework offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment. By offering specific solutions for each challenge, this framework aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
@@
  • The third framework presented in this SMS is the deployment challenge-solution framework for containerized applications in multi-cloud environment. This framework provides valuable insights and practical guidance to practitioners and researchers alike, aiding practitioners in navigating the complexities of multi-cloud deployment through specific solutions to challenges, while also serving as a comprehensive resource for researchers to delve into various aspects of containerized application deployment, orchestration, security, scalability, and more, thus contributing to the advancement of knowledge and innovation in the field of multi-cloud deployment.
@@
  • Fourth framework presented in this SMS is monitoring challenge-solution framework for containerized applications in multi-cloud environment. This framework serves as a comprehensive guide for practitioners, offering specific solutions to challenges related to monitoring performance, resource management, security, optimization, system complexity, and multi-cloud coordination, facilitating their understanding of complex monitoring scenarios and providing actionable insights. Furthermore, researchers can benefit from this framework by gaining a deeper insight into the nuances of monitoring containerized applications in multi-cloud environment, and it can serve as a foundation for further research and innovation in the field of multi-cloud monitoring.
@@
The findings of this SMS will benefit researchers who are interested in understanding the state of research on containerized applications in multi-cloud environment and conducting further investigations to address the open research issues highlighted in Section 4. Additionally, the insights gained from this SMS will support knowledge transfer to practitioners by providing insights into the challenges, solutions, and methods for monitoring, securing, and optimizing containerized applications in multi-cloud environment. We emphasize the importance for practitioners to develop targeted solutions that effectively tackle monitoring, security, and performance degradation concerns within multi-cloud environment deployment. As a future endeavor, we intend to enhance our SMS by conducting industrial case studies with companies, thereby obtaining practical insights and perspectives from practitioners on the effectiveness and applicability of our proposed frameworks in real-world scenarios. This approach will allow us to bridge the gap between research and practical implementations concerning containerized applications in multi-cloud environment and contribute to the advancement of both academic and industry understanding in this domain.
@@ Acknowledgments
This work is partially sponsored by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.
@@ 8. Appendices
List of selected studies for containerization in multi-cloud environments
@@
A component-based adaptation approach for multi-cloud applications
IEEE Conference on Computer Communications (INFOCOM) Workshops
2014 (Almeida et al., 2014)
S02
@@
A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications
International Conference on Web Information Systems Engineering (WISE)
2019 (Jha et al., 2019)
S03
@@ A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications
A framework to support multi-cloud collaboration
IEEE World Congress on Services (SERVICES)
2020 (Hua et al., 2020)
S04
@@
A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes
International Conference on Communications (ICC)
2021 (Ramos et al., 2021)
S05
@@ A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes
A multi-cloud service mesh approach applied to Internet of Things
Annual Conference of Industrial Electronics Society (IECON)
2022 (Gattobigio et al., 2022)
S06
@@
A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud
International Conference on Computer Supported Cooperative Work in Design (CSCWD)
2022 (Deng et al., 2022)
S07
@@ A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud
A Service Driver Based Application Execution and Development Method in Multi-cloud Context
International Conference on Data Science and Computer Application (ICDSCA)
2021 (Zhang et al., 2021)
S08
An adaptable framework to deploy complex applications onto multi-cloud platforms
@@ A Service Driver Based Application Execution and Development Method in Multi-cloud Context
International Conference on Computing Communication Technologies - Research, Innovation, and Vision for Future (RIVF)
2015 (Pham et al., 2015)
S09
An agile container-based approach to TaaS
@@
International Congress Future Telecommunications: Infrastructure and Sustainability (FITCE)
2016 (Verdugo et al., 2017)
S10
@@
An Efficient and Reliable Multi-Cloud Provider Monitoring Solution
Conference on Global Communications (GLOBECOM)
2022 (Sabbioni et al., 2020)
S11
@@
An Efficient Multi-Keyword Search Scheme over Encrypted Data in Multi-Cloud Environment
International Conference on Smart Cloud (SmartCloud)
2015 (He et al., 2022)
S12
@@
An Empirical Investigation of a Fault Tolerant Containerized Application Deployment
International Conference on Informatics (ICI)
2015 (Bisht and Kaur, 2022)
S13
@@
Applying Computational Intelligence for enhancing the dependability of multi-cloud systems using Docker Swarm
Symposium Series on Computational Intelligence (SSCI)
2017 (Naik, 2016a)
S14
@@
Automatically Design Heuristics for Multi-Objective Location-Aware Service Brokering in Multi-Cloud
International Conference on Services Computing (SCC)
2020 (Chen et al., 2022)
S15
@@
Autonomic fine-grained migration and replication of component-based applications across multi-clouds
National Foundation for Science and Technology Development Conference on Information and Computer Science (NICS)
2022 (Pham and Pham, 2015)
S16
@@
Band-Area Resource Management Platform and Accelerated Particle Swarm Optimization Alg. for Container Deployment
IEEE Access
2016 (Ouyang et al., 2022)
S17
@@
Building a Multi-tenant Cloud Service from Legacy Code with Docker Containers
International Conference on Cloud Engineering (IC2E)
2022 (Slominski et al., 2015)
S18
@@
Building cloud native application - Analysis for multi-component application deployment
International Conference on Computer Communication and Informatics (ICCCI)
2021 (Pai and Kumar, 2021)
S19
@@
CLAMS: Cross-layer Multi-cloud Application Monitoring-as-a-Service Framework
International Conference on Services Computing (SCC)
2014 (Alhamazani et al., 2014)
S20
@@
Cloud MF: Applying MDE to Tame the Complexity of Managing Multi-Cloud Applications
International Conference on Utility and Cloud Computing (UCC)
2014 (Ferry et al., 2014)
S21
@@
Cluster as a Service: A Container Based Cluster Sharing Approach with Multi-user Support
IEEE Symposium on Service-Oriented System Engineering (SOSE)
2016 (Cui et al., 2016)
S22
Container Deployment Strategy for Edge Networking
@@
Workshop on Middleware for Edge Clouds and Cloudlets (MECC)
2019 (Wong et al., 2019)
S23
@@
Container-Based Performance Isolation for Multi-Tenant SaaS Applications in Micro-Service Architecture
Journal of Physics: Conference Series
2020 (Wang et al., 2020)
S24
Containerization and the PaaS Cloud
IEEE Cloud Computing
2015 (Pahl, 2015)
S25
@@
ContainerLeaks: Emerging Security Threats of Information Leakages in Container Clouds
International Conference on Dependable Systems and Networks (DSN)
2017 (Gao et al., 2017)
S26
@@
ContMEC: An Architecture of Multi-access Edge Computing for Offloading Container-Based Mobile Applications
IEEE International Conference on Communications (ICC)
2022 (Watanabe et al., 2022)
S27
@@
DECIDE: An Extended DevOps Framework for Multi-Cloud Applications
International Conference on Cloud and Big Data Computing (ICCBDC)
2022 (Alonso et al., 2019)
S28
Delivering Elastic Containerized Cloud Applications to Enable DevOps
@@
International Workshop on Software Engineering for Adaptive and Self-Managing Systems (SEAMS)
2017 (Barna et al., 2017)
S29
Deployment of Multi-Tier Application on Cloud and Continuous Monitoring using Kubernetes
@@
International Conference on Electrical, Electronics, Communication, Computer Technologies and Optimization Techniques (ICEECCOT)
2021 (Gupta et al., 2021)
S30
@@
Deployment-Time Multi-Cloud Application Security
International Conference on Smart Computing (SMARTCOMP)
2017 (Sheridan et al., 2017)
S31
@@
DEVSML 3.0 Stack: Rapid deployment of DEVS Farm in distributed cloud environment using microservices and containers
Symposium on Theory of Modeling and Simulation (TMS/DEVS)
2017 (Mittal and Risco-Martín, 2017)
S32
@@
Disaster Recovery and Business Continuity for Database Services in Multi-Cloud
International Conference on Computer Applications & Information Security (ICCAIS)
2018 (Al-shammari and Alwan, 2018)
S33
@@
Dynamic On-Demand Fog Formation Offering On-the-Fly IoT Service Deployment
IEEE Transactions on Network and Service Management
2020 (Sami and Mourad, 2020)
S34
@@
Dynamic security assurance in multi-cloud DevOps
IEEE Conference on Communications and Network Security (CNS)
2017 (Rios et al., 2017)
S35
@@
Engineering and Experimentally Benchmarking a Container-based Edge Computing System
International Conference on Communications (ICC)
2020 (Carpio et al., 2020)
S36
@@
FogDocker: Start Container Now, Fetch Image Later
International Conference on Utility and Cloud Computing (UCC)
2019 (Lorenzo et al., 2019)
S37
@@
Importance of Application-Level Resource Management in Multi-Cloud Deployments
International Conference on Cloud Engineering (IC2E)
2019 (Dimitrijevic et al., 2019)
S38
@@
Knowledge-Engineered Multi-Cloud Resource Brokering for Application Workflow Optimization
IEEE Transactions on Network and Service Management
2022 (Pandey et al., 2022)
S39
@@
Label-affinity-Scheduler: Considering Business Requirements in Container Scheduling for Multi-Cloud and Multi-Tenant
Brazilian Symposium on Computing System Engineering (SBESC)
2022 (Altran et al., 2022)
S40
@@
Leveraging kernel security mechanisms to improve container security: A survey
International Conference on Availability, Reliability and Security (ARES)
2019 (Bélair et al., 2019)
S41
@@
Location-Aware and Budget-Constrained Service Deployment for Composite Applications in Multi-Cloud Environment
IEEE Transactions on Parallel and Distributed Systems
2020 (Shi et al., 2020)
S42
@@
MicroCloud: A Container-Based Solution for Efficient Resource Management in the Cloud
International Conference on Smart Cloud (SmartCloud)
2016 (Baresi et al., 2016)
S43
@@
Mobile Edge as Part of the Multi-Cloud Ecosystem: A Performance Study
Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP)
2019 (Dreibholz et al., 2019)
S44
@@
Modeling and Automated Deployment of Serverless Applications Using TOSCA
Conference on Service-Oriented Computing and Applications (SOCA)
2018 (Wurster et al., 2018)
S45
@@
Monitoring Elastically Adaptive Multi-Cloud Services
IEEE Transactions on Cloud Computing
2018 (Trihinas et al., 2015)
S46

### v2

@@ 6.2.2. Security of Containerized Clouds
Nicolae et al. (Casalicchio, 2019) analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework. Their research identifies potential attack scenarios and security enforcement mechanisms, aiming to enhance security guarantees for cloud operators in a multi-cloud setting. Mohammad et al. (Bokhari et al., 2018) offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models. They discuss the security requirements for public and private clouds across different service models, aiming to provide researchers with a comprehensive understanding of cloud computing’s potential and security challenges. Sari et al. (Sultan et al., 2019) survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape. Their analysis covers both software-based solutions utilizing Linux kernel features and hardware-based solutions for enhanced security. The study aims to clarify container security requirements and encourage further research in addressing potential vulnerabilities and attacks. Hendrik et al. (Graupner et al., 2015b) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective. They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control. Rohan et al. (Gupta et al., 2016) propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption. This approach aims to secure data while optimizing resource usage and cost, highlighting the benefits of container technology in improving cloud computing security and efficiency.
@@ 6.3. Conclusive Summary
Current survey-based studies and systematic reviews, discussed above (e.g., (Casalicchio and Iannucci, 2020) (Pahl et al., 2017)), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems. These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends. Table 16 presents an overview of the differences between this study and existing secondary studies. The symbols used in the table are as follows:  indicates “Yes”,  indicates “No”, and $\triangle$ indicates “Partially”. After thoroughly comparing our results with existing studies (e.g., (Pahl et al., 2017), (Casalicchio and Iannucci, 2020), (Netaji and Bhole, 2021)), it is clear that none of the results are fully identical with our findings. Our study distinguishes itself from prior research in several ways. First, our work systematically categorizes container utilization in multi-cloud environment into coherent categories and themes. Some of the key themes are Scalability and High Availability, Performance and Optimization, Security and Privacy, and Multi-Cloud Container Monitoring and Adaptation. We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided. Furthermore, we explored the QAs and associated tactics for containerization in multi-cloud environment. This contribution offers practitioners actionable insights into tailoring their containerization strategies to meet desired quality standards. Lastly, our study introduces four distinct frameworks consisting of challenges and solutions related to security, automation, deployment, and monitoring within the context of containerization in multi-cloud environment.
Table 16. Comparison of This SLR Results with Existing Secondary Studies. MS indicate “Mix Study”, LR indicate “Literature Review”  indicates “Yes”,  indicates “No,” and $\triangle$ indicates “Partially.”
Contributions This Study (Pahl et al., 2017) (Casalicchio and Iannucci, 2020) (Netaji and Bhole, 2021) (Bentaleb et al., 2022) (Watada et al., 2019) (Saxena et al., 2021) (Imran et al., 2020) (Casalicchio, 2019)
Number of selected studies 121 46 97 64 40 15 11
Study Types SMS SMS LR LR LR MS LR LR MS
Container Roles (Section 3.2.1) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Container Implementation Strategies (Section 3.2.2) $\triangle$ $\triangle$ $\triangle$
Pattern and Strategies (Section 3.3) $\triangle$ $\triangle$
Quality Attributes and Tactics (Section 3.4) $\triangle$ $\triangle$
Security Challenges and Solution (Section 3.5) $\triangle$ $\triangle$ $\triangle$
Automation Challenges and Solution (Section 3.6) $\triangle$ $\triangle$ $\triangle$
Deployment Challenges and Solution (Section 3.7) $\triangle$
Monitoring Challenges and Solution (Section 3.8) $\triangle$
Tools and Frameworks (Section 3.9) $\triangle$ $\triangle$ $\triangle$
@@ 7. Conclusions
This SMS presents the current state of research regarding containerization in multi-cloud environment, focusing on aims of selected studies, the roles of containers, containers implementation strategies, architectural patterns and strategies, quality attributes and corresponding tactics. Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment. From an initial pool of extensive search results, 86 relevant studies were finally chosen to answer the research questions presented in Table 1. The key findings of this SMS are listed below.
  • The findings of this SMS reveal various insights regarding SMS demographics, publishers, publication types, authors’ affiliations, and research themes. The yearly distribution reached its peak in 2016, signifying a shifting level of interest over time. The field of publishing is largely dominated by IEEE (55.37%), with conferences emerging as the primary avenue for publication. Notably, academia (75.02%) stands out as the leading affiliation among authors. The prominent research themes encompass Orchestration and Management, Scalability and High Availability, Performance and Optimization, and Security and Privacy. These themes underscore crucial aspects such as container management, scalability, performance, and security within the context of multi-cloud environment.
  • This SMS identifies and classifies a total of 98 patterns and strategies across 10 subcategories and 4 categories for container-based applications in multi-cloud environment. These patterns and strategies encompass a wide range of aspects such as cloud architecture models, communication and networking, deployment, service models, cloud management, container management, edge computing, IoT strategies, security, resilience, fault-tolerance, and cloud migration. These findings are significant for the application and dvancement of containerization-based applications in multi-cloud environment, providing valuable insights into optimizing architecture, communication, management, security, and migration strategies.
@@
  • This SMS presents a catalogs for the automation challenge-solution catalogs for containerized applications in multi-cloud environment. The automation challenge-solution catalog offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment. By offering specific solutions for each challenge, this catalog aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
@@
  • The third catalog presented in this SMS is the deployment challenge-solution catalog for containerized applications in multi-cloud environment. This catalog provides valuable insights and practical guidance to practitioners and researchers alike, aiding practitioners in navigating the complexities of multi-cloud deployment through specific solutions to challenges, while also serving as a comprehensive resource for researchers to delve into various aspects of containerized application deployment, orchestration, security, scalability, and more, thus contributing to the advancement of knowledge and innovation in the field of multi-cloud deployment.
@@
  • Fourth catalog presented in this SMS is monitoring challenge-solution catalog for containerized applications in multi-cloud environment. This catalog serves as a comprehensive guide for practitioners, offering specific solutions to challenges related to monitoring performance, resource management, security, optimization, system complexity, and multi-cloud coordination, facilitating their understanding of complex monitoring scenarios and providing actionable insights. Furthermore, researchers can benefit from this catalog by gaining a deeper insight into the nuances of monitoring containerized applications in multi-cloud environment, and it can serve as a foundation for further research and innovation in the field of multi-cloud monitoring.
@@
The findings of this SMS will benefit researchers who are interested in understanding the state of research on containerized applications in multi-cloud environment and conducting further investigations to address the open research issues highlighted in Section 4. Additionally, the insights gained from this SMS will support knowledge transfer to practitioners by providing insights into the challenges, solutions, and methods for monitoring, securing, and optimizing containerized applications in multi-cloud environment. We emphasize the importance for practitioners to develop targeted solutions that effectively tackle monitoring, security, and performance degradation concerns within multi-cloud environment deployment. As a future endeavor, we intend to enhance our SMS by conducting industrial case studies with companies, thereby obtaining practical insights and perspectives from practitioners on the effectiveness and applicability of our proposed catalogs. This approach will allow us to bridge the gap between research and practical implementations concerning containerized applications in multi-cloud environment and contribute to the advancement of both academic and industry understanding in this domain.
Data Availability
The dataset of this study has been provided online: https://zenodo.org/record/10732611.
@@ Acknowledgments
This research is funded by Business Finland through QLEAP (2022-24) project, and by the NSFC China under Grant No. 62172311.
@@ 8. Appendices
{xltabular}
—c—p4.5cm—p4.5cm—c—c— List of selected studies for containerization in multi-cloud environments
@@
S01 A component-based adaptation approach for multi-cloud applications IEEE Conference on Computer Communications (INFOCOM) Workshops 2014 (Almeida et al., 2014)
@@
S02 A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications International Conference on Web Information Systems Engineering (WISE) 2019 (Jha et al., 2019)
@@
S03 A framework to support multi-cloud collaboration IEEE World Congress on Services (SERVICES) 2020 (Hua et al., 2020)
@@
S04 A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes International Conference on Communications (ICC) 2021 (Ramos et al., 2021)
@@
S05 A multi-cloud service mesh approach applied to Internet of Things Annual Conference of Industrial Electronics Society (IECON) 2022 (Gattobigio et al., 2022)
@@
S06 A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud International Conference on Computer Supported Cooperative Work in Design (CSCWD) 2022 (Deng et al., 2022)
@@
S07 A Service Driver Based Application Execution and Development Method in Multi-cloud Context International Conference on Data Science and Computer Application (ICDSCA) 2021 (Zhang et al., 2021)
@@
S08 An adaptable framework to deploy complex applications onto multi-cloud platforms International Conference on Computing Communication Technologies - Research, Innovation, and Vision for Future (RIVF) 2015 (Pham et al., 2015)
@@
S09 An agile container-based approach to TaaS International Congress Future Telecommunications: Infrastructure and Sustainability (FITCE) 2016 (Verdugo et al., 2017)
@@
S10 An Efficient and Reliable Multi-Cloud Provider Monitoring Solution Conference on Global Communications (GLOBECOM) 2022 (Sabbioni et al., 2020)
@@
S11 An Efficient Multi-Keyword Search Scheme over Encrypted Data in Multi-Cloud Environment International Conference on Smart Cloud (SmartCloud) 2015 (He et al., 2022)
@@
S12 An Empirical Investigation of a Fault Tolerant Containerized Application Deployment International Conference on Informatics (ICI) 2015 (Bisht and Kaur, 2022)
@@
S13 Applying Computational Intelligence for enhancing the dependability of multi-cloud systems using Docker Swarm Symposium Series on Computational Intelligence (SSCI) 2017 (Naik, 2016a)
@@
S14 Automatically Design Heuristics for Multi-Objective Location-Aware Service Brokering in Multi-Cloud International Conference on Services Computing (SCC) 2020 (Chen et al., 2022)
@@
S15 Autonomic fine-grained migration and replication of component-based applications across multi-clouds National Foundation for Science and Technology Development Conference on Information and Computer Science (NICS) 2022 (Pham and Pham, 2015)
@@
S16 Band-Area Resource Management Platform and Accelerated Particle Swarm Optimization Alg. for Container Deployment IEEE Access 2016 (Ouyang et al., 2022)
@@
S17 Building a Multi-tenant Cloud Service from Legacy Code with Docker Containers International Conference on Cloud Engineering (IC2E) 2022 (Slominski et al., 2015)
@@
S18 Building cloud native application - Analysis for multi-component application deployment International Conference on Computer Communication and Informatics (ICCCI) 2021 (Pai and Kumar, 2021)
@@
S19 CLAMS: Cross-layer Multi-cloud Application Monitoring-as-a-Service Framework International Conference on Services Computing (SCC) 2014 (Alhamazani et al., 2014)
@@
S20 Cloud MF: Applying MDE to Tame the Complexity of Managing Multi-Cloud Applications International Conference on Utility and Cloud Computing (UCC) 2014 (Ferry et al., 2014)
@@
S21 Cluster as a Service: A Container Based Cluster Sharing Approach with Multi-user Support IEEE Symposium on Service-Oriented System Engineering (SOSE) 2016 (Cui et al., 2016)
@@
S22 Container Deployment Strategy for Edge Networking Workshop on Middleware for Edge Clouds and Cloudlets (MECC) 2019 (Wong et al., 2019)
@@
S23 Container-Based Performance Isolation for Multi-Tenant SaaS Applications in Micro-Service Architecture Journal of Physics: Conference Series 2020 (Wang et al., 2020)
S24 Containerization and the PaaS Cloud IEEE Cloud Computing 2015 (Pahl, 2015)
@@
S25 ContainerLeaks: Emerging Security Threats of Information Leakages in Container Clouds International Conference on Dependable Systems and Networks (DSN) 2017 (Gao et al., 2017)
@@
S26 ContMEC: An Architecture of Multi-access Edge Computing for Offloading Container-Based Mobile Applications IEEE International Conference on Communications (ICC) 2022 (Watanabe et al., 2022)
@@
S27 DECIDE: An Extended DevOps Framework for Multi-Cloud Applications International Conference on Cloud and Big Data Computing (ICCBDC) 2022 (Alonso et al., 2019)
@@
S28 Delivering Elastic Containerized Cloud Applications to Enable DevOps International Workshop on Software Engineering for Adaptive and Self-Managing Systems (SEAMS) 2017 (Barna et al., 2017)
@@
S29 Deployment of Multi-Tier Application on Cloud and Continuous Monitoring using Kubernetes International Conference on Electrical, Electronics, Communication, Computer Technologies and Optimization Techniques (ICEECCOT) 2021 (Gupta et al., 2021)
@@
S30 Deployment-Time Multi-Cloud Application Security International Conference on Smart Computing (SMARTCOMP) 2017 (Sheridan et al., 2017)
@@
S31 DEVSML 3.0 Stack: Rapid deployment of DEVS Farm in distributed cloud environment using microservices and containers Symposium on Theory of Modeling and Simulation (TMS/DEVS) 2017 (Mittal and Risco-Martín, 2017)
@@
S32 Disaster Recovery and Business Continuity for Database Services in Multi-Cloud International Conference on Computer Applications & Information Security (ICCAIS) 2018 (Al-shammari and Alwan, 2018)
@@
S33 Dynamic On-Demand Fog Formation Offering On-the-Fly IoT Service Deployment IEEE Transactions on Network and Service Management 2020 (Sami and Mourad, 2020)
@@
S34 Dynamic security assurance in multi-cloud DevOps IEEE Conference on Communications and Network Security (CNS) 2017 (Rios et al., 2017)
@@
S35 Engineering and Experimentally Benchmarking a Container-based Edge Computing System International Conference on Communications (ICC) 2020 (Carpio et al., 2020)
@@
S36 FogDocker: Start Container Now, Fetch Image Later International Conference on Utility and Cloud Computing (UCC) 2019 (Lorenzo et al., 2019)
@@
S37 Importance of Application-Level Resource Management in Multi-Cloud Deployments International Conference on Cloud Engineering (IC2E) 2019 (Dimitrijevic et al., 2019)
@@
S38 Knowledge-Engineered Multi-Cloud Resource Brokering for Application Workflow Optimization IEEE Transactions on Network and Service Management 2022 (Pandey et al., 2022)
@@
S39 Label-affinity-Scheduler: Considering Business Requirements in Container Scheduling for Multi-Cloud and Multi-Tenant Brazilian Symposium on Computing System Engineering (SBESC) 2022 (Altran et al., 2022)
@@
S40 Leveraging kernel security mechanisms to improve container security: A survey International Conference on Availability, Reliability and Security (ARES) 2019 (Bélair et al., 2019)
@@
S41 Location-Aware and Budget-Constrained Service Deployment for Composite Applications in Multi-Cloud Environment IEEE Transactions on Parallel and Distributed Systems 2020 (Shi et al., 2020)
@@
S42 MicroCloud: A Container-Based Solution for Efficient Resource Management in the Cloud International Conference on Smart Cloud (SmartCloud) 2016 (Baresi et al., 2016)
@@
S43 Mobile Edge as Part of the Multi-Cloud Ecosystem: A Performance Study Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP) 2019 (Dreibholz et al., 2019)
@@
S44 Modeling and Automated Deployment of Serverless Applications Using TOSCA Conference on Service-Oriented Computing and Applications (SOCA) 2018 (Wurster et al., 2018)
@@
S45 Monitoring Elastically Adaptive Multi-Cloud Services IEEE Transactions on Cloud Computing 2018 (Trihinas et al., 2015)